Ordered pairs?
X-y= 11. 3x+10y=-6

Answer by mananth(16946)

1 x -1 y = 11 .............1

3 x + 10 y = -6 .............2
Eliminate y
multiply (1)by 10
Multiply (2) by 1
10 x -10 y = 110
3 x + 10 y = -6
Add the two equations
13 x = 104
/ 13
x = 8
plug value of x in (1)
1 x -1 y = 11
8 -1 y = 11
-1 y = 11 -8
-1 y = 3
y = -3
(8,-3)